OpenStudy (anonymous):

Zach wrote the formula w(w – 1)(5w + 2) for the volume of a rectangular prism he is designing, with width w, which is always has a positive value greater than 1. Find the product and then classify this polynomial by degree and by number of terms.

OpenStudy (anonymous):

You have to foil. After you foiled the equations the highest number exponent will be the degree of the equation. The number of terms is simply how many things you have at the end, i.e, 4x^2+2x+1 has 3 terms, 4x^2, 2x, and 1 are all terms.

OpenStudy (asapbleh):

\[w \left( w -1 \right)\left( 5w +2 \right)\] First you will need to multiply w(w-1) w*w= \[w ^{2}\] w*-1=-w so now its \[(w ^{2}-w)(5w+2)\]

OpenStudy (asapbleh):

Try to multiply the outermost part, which is w^2*5w*, then w^2*2, then -w*5w, and finally -w*2. The degree will be to the 3rd.

OpenStudy (anonymous):

5w^5-3w^4-2w^3 correct @asapbleh

OpenStudy (anonymous):

?

OpenStudy (asapbleh):

no that doesnt seem about right,  try again